# Towards shutdownable agents via stochastic choice

### v1

@@ Appendix B Proof
    For all other trajectory-lengths $a_{r}$, $p_{\pi}(a_{r})=p_{\pi^{\prime}}(a_{r})$.
The expected return of $\pi^{\prime}$ is greater than the expected return of $\pi$.
In other words, we can increase the expected return of any maximally USEFUL policy by shifting probability mass away from trajectory-length $a_{i}$ and towards trajectory-length $a_{j}$, up until the point where $p(a_{i})=p(a_{j})$.
To prove Lemma 1, it’s convenient to suppose (contrary to fact) that the agent can distinguish between different mini-episodes in the meta-episode and hence can select different probability distributions over trajectory-lengths in different mini-episodes. We make this supposition in Lemma 2 and then use Lemma 2 to prove Lemma 1.
Lemma 2.
For any maximally USEFUL policy $\pi$ and any pair of trajectory-lengths $a_{i}$ and $a_{j}$ such that:
  1. 1.
    For some number $e_{1}>0$ of mini-episodes in the meta-episode, $p_{\pi}(a_{i})>p_{\pi}(a_{j})$ (with $\mu$ denoting the sum of $p_{\pi}(a_{i})$ and $p_{\pi}(a_{j})$).
  2. 2.
    For some number $e_{2}\geq 0$ of mini-episodes in the meta-episode, $p_{\pi}(a_{i})=p_{\pi}(a_{j})=\frac{\mu}{2}$.
  3. 3.
    $e_{1}+e_{2}=N-1$, where $N$ is the number of mini-episodes in the meta-episode (i.e. there’s only one remaining mini-episode with the relative sizes of $p_{\pi}(a_{i})$ and $p_{\pi}(a_{j})$ undetermined by conditions 1 and 2 above).
Then the expected return of $\pi$ would be greater if $p_{\pi}(a_{i})=p_{\pi}(a_{j})=\frac{\mu}{2}$ in the remaining mini-episode than if $p_{\pi}(a_{i})>p_{\pi}(a_{j})$ (with $p_{\pi}(a_{i})+p_{\pi}(a_{j})=\mu$) in the remaining mini-episode.
In other words, if a maximally USEFUL policy $\pi$ is biased towards some trajectory-length $a_{i}$ over some other trajectory-length $a_{j}$ in some set of mini-episodes (and isn’t biased towards $a_{j}$ over $a_{i}$ in any mini-episodes), then we can increase $\pi$’s expected return by ensuring that $p_{\pi}(a_{i})=p_{\pi}(a_{j})=\frac{\mu}{2}$ (rather than $p_{\pi}(a_{i})>p_{\pi}(a_{j})$ with $p_{\pi}(a_{i})+p_{\pi}(a_{j})=\mu$) in the remaining mini-episode.
Here’s the proof of Lemma 2. Given antecedent conditions 1-3 of Lemma 2 above, $a_{i}$’s probability distribution over times-chosen in the first $N-1$ mini-episodes stochastically dominates $a_{j}$’s probability distribution over times-chosen in the first $N-1$ mini-episodes. That is to say:
  1. 1.
    For each possible number of times-chosen $w$, $a_{i}$’s probability of being chosen at least as many times as $w$ is at least as great as $a_{j}$’s probability of being chosen at least as many times as $w$.
  2. 2.
    For some possible number of times-chosen $w$, $a_{i}$’s probability of being chosen at least as many times as $w$ is greater than $a_{j}$’s probability of being chosen at least as many times as $w$.
Now recall that, for any trajectory-lengths $a_{i}$ and $a_{j}$, if $a_{i}$ has so far been chosen more times in the meta-episode than $a_{j}$, the return for choosing $a_{j}$ in the next mini-episode is greater than the return for choosing $a_{i}$ in the next mini-episode.
This fact (in combination with the fact that $a_{i}$’s probability distribution over times-chosen in the first $N-1$ mini-episodes stochastically dominates $a_{j}$’s probability distribution over times-chosen in the first $N-1$ mini-episodes) implies that $a_{j}$’s probability distribution over returns in the remaining mini-episode stochastically dominates $a_{i}$’s probability distribution over returns in the remaining mini-episode. That is to say:
  1. 1.
    For each possible return in the remaining mini-episode $r$, $a_{j}$’s probability of delivering at least $r$ is at least as great as $a_{i}$’s probability of delivering at least $r$.
  2. 2.
    For some possible return in the remaining mini-episode $r$, $a_{j}$’s probability of delivering at least $r$ is greater than $a_{i}$’s probability of delivering at least $r$.
That implies that the expected return of choosing $a_{j}$ in the remaining mini-episode is greater than the expected return of choosing $a_{i}$ in the remaining mini-episode. That in turn implies that the expected return of choosing $a_{i}$ and $a_{j}$ each with probability $\frac{\mu}{2}$ is greater than the expected return of choosing $a_{i}$ with probability greater than $a_{j}$ (given that $p(a_{i})+p(a_{j})=\mu$). That in turn implies that the expected return of $\pi$ across the meta-episode would be greater if $a_{i}$ and $a_{j}$ were each chosen with probability $\frac{\mu}{2}$ in the remaining mini-episode (as opposed to choosing $a_{i}$ with probability greater than $a_{j}$).
That completes the proof of Lemma 2. Now we use it to prove Lemma 1. Let $\pi$ be a policy such that $p_{\pi}(a_{i})>p_{\pi}(a_{j})$ in each mini-episode. As above, let $\mu$ denote the sum of $p_{\pi}(a_{i})$ and $p_{\pi}(a_{j})$. By Lemma 2, we’d increase expected return across the meta-episode if the policy were such that $p_{\pi}(a_{i})=p_{\pi}(a_{j})=\frac{\mu}{2}$ in the last mini-episode. And then (also by Lemma 2), we’d increase expected return further if the policy were also such that $p_{\pi}(a_{i})=p_{\pi}(a_{j})=\frac{\mu}{2}$ in the second-to-last mini-episode, and so on all the way back to the second mini-episode.
Now suppose that $p_{\pi}(a_{i})=p_{\pi}(a_{j})=\frac{\mu}{2}$ in all mini-episodes except the first. We’re now trying to decide how to split probability mass between $a_{i}$ and $a_{j}$ in the first mini-episode. Given our reasoning above, $a_{i}$’s probability distribution over times-chosen in the last $N-1$ mini-episodes is identical to $a_{j}$’s probability distribution over times-chosen in the last $N-1$ mini-episodes, so their probability distributions over returns in the first mini-episode are identical too. Thus, their expected returns in the first mini-episode are also identical. Hence, expected return across the meta-episode is the same regardless of how we split $\mu$ between $p_{\pi}(a_{i})$ and $p_{\pi}(a_{j})$ in the first mini-episode.
Thus, if we start with a $\pi$ such that $p_{\pi}(a_{i})>p_{\pi}(a_{j})$ in each mini-episode, we strictly increase expected return by equalizing $p_{\pi}(a_{i})$ and $p_{\pi}(a_{j})$ in all mini-episodes except one, and equalizing $p_{\pi}(a_{i})$ and $p_{\pi}(a_{j})$ in the remaining mini-episode doesn’t decrease expected return. Hence, in any meta-episode containing more than one mini-episode, the expected return of $\pi^{\prime}$ (in which $p_{\pi^{\prime}}(a_{i})=p_{\pi^{\prime}}(a_{j})=\frac{\mu}{2}$ in each mini-episode) is greater than that of $\pi$ (in which $p_{\pi}(a_{i})>p_{\pi}(a_{j})$ with $p_{\pi}(a_{i})+p_{\pi}(a_{j})=\mu$ in each mini-episode). That completes the proof of Lemma 1.
@@
Now we use Lemma 1. Take any maximally USEFUL policy $\pi$. By Lemma 1, if there are any trajectory-lengths $a_{i}$ and $a_{j}$ such that $p(a_{i})>p(a_{j})$, then we can increase $\pi$’s expected return by equalizing $p(a_{i})$ and $p(a_{j})$. If we do so for each $a_{i}$ and $a_{j}$ such that $p(a_{i})>p(a_{j})$, we get $p(a_{i})=\frac{1}{k}$ for each of the $k$ available trajectory-lengths $a_{i}$. We then maximize:
$-\sum_{i=1}^{k}p(a_{i})\log_{2}p(a_{i})$
Therefore, optimal polices are maximally NEUTRAL.
We have thus proved that optimal policies for our DREST reward function are maximally USEFUL and maximally NEUTRAL.

### v2

@@ Appendix D Proof
    And for all other trajectory-lengths $l$, $Pr_{\pi}\{L=l\}=Pr_{\pi^{\prime}}\{L=l\}$,
Then $\mathbb{E}_{\pi^{\prime},E}(R)>\mathbb{E}_{\pi,E}(R)$.
Proof. Let $E$ be a meta-episode consisting of $n$ mini-episodes with $n>1$. Assume that each policy $\pi$ below is maximally USEFUL. Recall that $N_{e_{i}}(L=l)$ denotes the number of times that trajectory-length $l$ has been chosen prior to mini-episode $e_{i}$.
Note that the expected return of a policy $\pi$ in a meta-episode $e_{s}$ conditional on selecting a trajectory-length $x$ can be expressed as follows:
$\mathbb{E}_{\pi,e_{s}}(R|L=x)=\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)\\ +\sum_{i=1}^{s-1}\big{(}\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1-i)-\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-i)\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}$ (1)
Here is how to interpret this equation. Selecting trajectory-length $x$ in mini-episode $e_{s}$ is guaranteed to yield at least $\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)$: the expected return that would be had if $x$ were selected in all $s-1$ previous mini-episodes. In addition, there is a probability of $Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-2\}$ that selecting $x$ in $e_{s}$ yields $\big{(}\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-2)-\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)\big{)}$: the extra expected return that would be had if $x$ were selected in only $s-2$ previous mini-episodes. In addition, there is a probability of $Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-3\}$ that selecting $x$ in $e_{s}$ yields $\big{(}\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-3)-\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-2)\big{)}$: the extra expected return that would be had if $x$ were selected in only $s-3$ previous mini-episodes. And so on.
If policy $\pi$ is maximally USEFUL, then the expected return for selecting trajectory-length $x$ in mini-episode $e_{s}$ given that trajectory-length $x$ has been selected $b$ times prior to $e_{s}$ is:
$\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=b)=\lambda^{b-\frac{s-1}{k}}$
Therefore, the expected return of a policy $\pi$ in a meta-episode $e_{s}$ conditional on selecting a trajectory-length $x$ can be expressed as follows:
$\mathbb{E}_{\pi,e_{s}}(R|L=x)=\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}$ (2)
Similarly, the expected return of a policy $\pi$ in a meta-episode $e_{s}$ conditional on selecting a trajectory-length $y$ can be expressed as follows:
$\mathbb{E}_{\pi,e_{s}}(R|L=y)=\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\cdot Pr_{\pi}\{N_{e_{s}}(L=y)\leq s-1-i\}$ (3)
Therefore, the expected return of a policy $\pi$ in a meta-episode $e_{s}$ conditional on selecting either trajectory-length $x$ or trajectory-length $y$ can be expressed as follows:
$\mathbb{E}_{\pi,e_{s}}(R|L=x\lor L=y)=\\ Pr_{\pi,e_{s}}\{L=x\}\cdot\bigg{(}\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}\bigg{)}\\ +Pr_{\pi,e_{s}}\{L=y\}\cdot\bigg{(}\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\cdot Pr_{\pi}\{N_{e_{s}}(L=y)\leq s-1-i\}\bigg{)}$ (4)
Let $\pi_{n}$ be a policy that selects trajectory-length $x$ with greater probability than trajectory-length $y$ in each mini-episode $e_{1}$ to $e_{n}$ (denoted $e_{1}-e_{n}$). More precisely, $\pi_{n}$ is such that, for trajectory-lengths $x$ and $y$, $Pr_{\pi_{n},e_{1}-e_{n}}\{L=x\}>Pr_{\pi_{n},e_{1}-e_{n}}\{L=y\}$.
Let $Pr_{\pi_{n},e_{1}-e_{n}}\{L=x\}=\mu+\Delta$ and $Pr_{\pi_{n},e_{1}-e_{n}}\{L=y\}=\mu-\Delta$.
Let $\pi_{n-1}$ be identical to $\pi_{n}$ except that $\pi_{n-1}$ selects trajectory-lengths $x$ and $y$ with equal probability $\mu$ in the final mini-episode $e_{n}$. More precisely, $\pi_{n-1}$ is such that $Pr_{\pi_{n-1},e_{n}}\{L=x\}=Pr_{\pi_{n-1},e_{n}}\{L=y\}=\mu$. For all other trajectory-lengths $l$ besides $x$ and $y$, $Pr_{\pi_{n-1},e_{1}-e_{n}}\{L=l\}=Pr_{\pi_{n},e_{1}-e_{n}}\{L=l\}$.
(Note that $\pi_{n-1}$ implies one probability distribution over trajectory-lengths in the first $n-1$ mini-episodes $e_{1}$ to $e_{n-1}$ and implies a different probability distribution over trajectory-lengths in the final mini-episode $e_{n}$. Given that the environments in mini-episodes $e_{1}$ to $e_{n}$ are observationally-equivalent, policies like $\pi_{n-1}$ cannot be implemented. Nevertheless, it is useful to refer to policies like $\pi_{n-1}$ in proving Lemma D.2.)
Let $\pi_{n-2}$ be identical to $\pi_{n}$ except that $\pi_{n-2}$ selects trajectory-lengths $x$ and $y$ with the same probability $\mu$ in the final two mini-episodes $e_{n-1}$ to $e_{n}$. More precisely, $\pi_{n-2}$ is such that $Pr_{\pi_{n-2},e_{n-1}-e_{n}}\{L=x\}=Pr_{\pi_{n-2},e_{n-1}-e_{n}}\{L=y\}=\mu$.
And so on.
Let $\pi_{1}$ be identical to $\pi_{n}$ except that $\pi_{1}$ selects trajectory-lengths $x$ and $y$ with the same probability $\mu$ in all but the first mini-episode $e_{1}$. More precisely, $\pi_{1}$ is such that $Pr_{\pi_{1},e_{2}-e_{n}}\{L=x\}=Pr_{\pi_{1},e_{2}-e_{n}}\{L=y\}=\mu$.
Let $\pi_{0}$ be identical to $\pi_{n}$ except that $\pi_{0}$ selects trajectory-lengths $x$ and $y$ with the same probability $\mu$i n all mini-episodes $e_{1}$ to $e_{n}$. More precisely, $\pi_{0}$ is such that $Pr_{\pi_{0},e_{1}-e_{n}}\{L=x\}=Pr_{\pi_{0},e_{1}-e_{n}}\{L=y\}=\mu$.
We will prove that $\mathbb{E}_{\pi_{n},E}(R)<\mathbb{E}_{\pi_{0},E}(R)$. We will thereby prove Lemma D.2.
Consider a pair of policies $\pi_{a}$ and $\pi_{a-1}$ with $1\leq a\leq n$. We can express as follows the expected return of $\pi_{a-1}$ across the meta-episode $E$ conditional on selecting trajectory-length $x$ or $y$ in each mini-episode:
$\mathbb{E}_{\pi_{a-1},E}(R|L=x\lor L=y)=\mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (5)
The first term on the right-hand side is the expected return of $\pi_{a-1}$ in mini-episodes $e_{1}$ to $e_{a-1}$ conditional on selecting trajectory-length $x$ or $y$ in each of these mini-episodes. The middle two terms give the expected return of $\pi_{a-1}$ conditional on selecting trajectory-length $x$ or $y$ in mini-episode $e_{a}$: the first mini-episode in which $\pi_{a-1}$ selects trajectory-lengths $x$ and $y$ with equal probability $\mu$. The final term is the sum of expected returns of $\pi_{a-1}$ in the remaining mini-episodes conditional on selecting trajectory-length $x$ or $y$ in each of these mini-episodes.
Similarly, we can express as follows the expected return of $\pi_{a}$ across the meta-episode $E$ conditional on selecting trajectory-length $x$ or $y$ in each mini-episode:
$\mathbb{E}_{\pi_{a},E}(R|L=x\lor L=y)=\mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +(\mu+\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +(\mu-\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (6)
As above, the first term on the right-hand side is the expected return of $\pi_{a}$ in mini-episodes $e_{1}$ to $e_{a-1}$ conditional on selecting trajectory-length $x$ or $y$ in each of these mini-episodes. The middle two terms give the expected return of $\pi_{a}$ conditional on selecting trajectory-length $x$ or $y$ in mini-episode $e_{a}$: the last mini-episode in which $\pi_{a}$ selects trajectory-length $x$ with probability $\mu+\Delta$ and selects trajectory-length $y$ with probability $\mu-\Delta$. The final term is the sum of expected returns of $\pi_{a}$ in the remaining mini-episodes conditional on selecting trajectory-length $x$ or $y$ in each of these mini-episodes.
We now prove that $\pi_{a-1}$ has greater expected return than $\pi_{a}$. Since $\pi_{a-1}$ and $\pi_{a}$ are each maximally USEFUL, and since for all trajectory-lengths $l$ besides $x$ and $y$, $Pr_{\pi_{a-1},e_{1}-e_{n}}\{L=l\}=Pr_{\pi_{a},e_{1}-e_{n}}\{L=l\}$, we need only prove that $\mathbb{E}_{\pi_{a-1},E}(R|L=x\lor L=y)>\mathbb{E}_{\pi_{a},E}(R|L=x\lor L=y)$.
The statement to be proved can be expressed as follows:
$\mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\\ >\mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +(\mu+\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +(\mu-\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (7)
Since $\pi_{a-1}$ and $\pi_{a}$ are each maximally USEFUL, and since $Pr_{\pi_{a-1},e_{1}-e_{a-1}}\{L=x\}=Pr_{\pi_{a},e_{1}-e_{a-1}}\{L=x\}=\mu+\Delta$ and $Pr_{\pi_{a-1},e_{1}-e_{a-1}}\{L=x\}=Pr_{\pi_{a},e_{1}-e_{a-1}}\{L=x\}=\mu-\Delta$, it follows that $\mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)=\mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)$. We can thus cancel the first term on each side of the inequality. And then by simple algebra the inequality can be expressed as follows:
$\Delta\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}-Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\})\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}+Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}-Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\})\bigg{)}\bigg{)}>0$ (8)
By stipulation, $\Delta>0$. And since $10<\lambda<$ 1, $\lambda^{a-1-\frac{a-1}{k}}>0$ and $\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}>0$ for all $a$, $n$, and $k$. And since $Pr_{\pi_{a},e_{1}-e_{a}}\{L=x\}>Pr_{\pi_{a},e_{1}-e_{a}}\{L=y\}$, $Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}--Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\geq 0$ for all $a$ and $i$ and $Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}--Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}>0$ for all $a$ and some $i$ such that $11\leq i\leq a-$ 1. Therefore, the first term of the left-hand side above is strictly greater than zero.
And since, $\mu>0$, $\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}>0$ for all $j$, $i$, and $k$, and in each mini-episode $e_{s}$, $Pr_{\pi_{a-1},e_{s}}(L=x\lor L=y\}=Pr_{\pi_{a},e_{s}}(L=x\lor L=y\}=2\mu$, it follows that for all $a$, $n$, $\mu>0$, $k$:
$\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}+Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}-Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\})\bigg{)}\bigg{)}\geq 0$ (9)
Therefore, the left-hand side is strictly greater than zero. Therefore, $\mathbb{E}_{\pi_{a-1},E}(R|L=x\lor L=y)>\mathbb{E}_{\pi_{a},E}(R|L=x\lor L=y)$. Therefore, $\mathbb{E}_{\pi_{a-1},E}(R)>\mathbb{E}_{\pi_{a},E}(R)$. Therefore, $\mathbb{E}_{\pi_{0},E}(R)>\mathbb{E}_{\pi_{n},E}(R)$. That concludes the proof of Lemma D.2.
@@
Now we use Lemma D.2. For any maximally USEFUL policy $\pi$, if there are any trajectory-lengths $x$ and $y$ such that $Pr_{\pi,e_{1}-e_{n}}\{L=x\}>Pr_{\pi,e_{1}-e_{n}}\{L=y\}$, then the policy $\pi^{\prime}$ that is identical except that $Pr_{\pi^{\prime},e_{1}-e_{n}}\{L=x\}=Pr_{\pi^{\prime},e_{1}-e_{n}}\{L=y\}$ has greater expected return. So any policy $\pi^{*}$ that maximizes expected return must be such that, for any trajectory-lengths $x$ and $y$, $Pr_{\pi^{*},e_{1}-e_{n}}\{L=x\}=Pr_{\pi^{*},e_{1}-e_{n}}\{L=y\}$. Therefore, any policy $\pi^{*}$ that maximizes expected return must be maximally NEUTRAL. ∎